where a is the altitude above sea level (in miles). If a city has an atmospheric pressure of 11.19 psi, what is its altitude? (Recall that 1 mi = 5,280 ft. Round your answer to the nearest foot.)

Answer by Theo(13342)

you need the formula for atmospheric pressure in terms of altitude.

the formula is normally shown as P = 14.7 * e^(-.21*a).

if your formula is different because they take 14.7 out to more decimal places, then you need to use what you are given instead.

P is the atmospheric pressure.
a is the altitude in miles.

in your problem, P is given as 11.19 and you want to solve for a.

the equation becomes:

11.19 = 14.7 * e^(-.21*a)

divide both sides of this equation by 14.7 to get:

11.19 / 14.7 = e^(-.21*a)

take the natural log of both sides of this equation to get:

ln(11.19/14.7) = ln(e^(-.21*a)

since ln(e^(-.21*a) is equal to -.21*a*ln(e), your equation becomes:

ln(11.19/14.7) = -.21*a*ln(e)

since ln(e) = 1, your equation becomes:

ln(11.19/14.7) = -.21*a

divide both sides of this equation by -.21 to get:

ln(11.19/14.7)/-.21 = a

solve for a to get:

a = 1.299176... miles

since there are 5280 feet in a mile, then you get:

a = 6859.6495... feet which you can round to 6860 feet.